A meterologist reports that the probability that it will rain on Saturday is 30% and the probability that it will rain on Sunday is 40%. If the two events are independent, what is the probability that it will rain on at least one of the two days?
Answer #1

Solution:

independent events def:

if Both A and B are indepndent the probability of both occuring is given by

A\bigcap B=A.Bp(A\bigcap B)=p(A).p(B)

Here A=saturday

B=sunday

=0.3*0.4=0.12

the probability that it will rain on at least one of the two days?

=P(AUB)

=P(A)+P(B)-P(A and B)

=0.3+0.4-0.12

=0.7-0.12

=0.58

ANSWER:0.58